Humans carry the dominant allele T can taste the bitter substance PTC, while tt homozygotes cannot. In a population where the frequency of the T allele is 0.4, what is the probability that a particular taster is a heterozygote?

Please explain the answer.
Answer #1

Given that,

Frequency of Taster allele, T = Fr(T) = p = 0.4
q = 1 - p = 1 - 0.4 = 0.6

Thus the frequency of the non-taster allele, t, in the population = q = 0.6

Therefore, frequency of Heterozygotes = 2pq = 2 * p * q = 2 * 0.4 * 0.6 = 0.48

Since both, Homozygous Dominant individuals and Heterozygous individuals can taste PTC:

Frequency of Tasters = p2 + 2pq = 0.4 * 0.4 + 0.48 = 0.16 + 0.48 = 0.64

Probability that a taster is a heterozygote = Frequency of Heterozygotes / Frequency of Tasters

= 0.48/0.64
= 0.75

The probability that a taster is Hetrozygous is 0.75 or 75%